# REAL-Prover: Retrieval Augmented Lean Prover for Mathematical Reasoning

### v2

@@ 3.2 Benchmark
As our prover highlights its capabilities on college-level mathematics, we created our own benchmark, FATE-M, which focuses specifically on algebraic formal statements to test model’s performance on rich algebra structures. In addition, we also follow prior work xin2024deepseekprover ; theoremllama ; li2024hunyuanprover ; lin2025goedel to evaluate our prover’s general capabilities using the ProofNet and MiniF2F benchmarks.
@@
The benchmark was created by extracting problems from 12 university-level abstract algebra textbooks (for a full list, see Appendix B), which were then formalized by students. Each formalized statement was rigorously verified by Mathlib contributors and PhD students in algebra to ensure it accurately represents the original mathematical meaning. Additionally, Lean-checked proofs are provided for all statements to guarantee their correctness.
@@
example {G : Type*} {p : ℕ} [Group G] [Fintype G] (pp : p.Prime) (ord : Fintype.card G = p ^ 2) : P : Subgroup G, (P.Normal) (Fintype.card P = p) :=
@@ 4.1 ProofNet Result
Goedel Prover lin2025goedel 32 15.6 %
DeepSeek-Prover-V1.5-SFT xin2024deepseekprover 128 15.9 %
DeepSeek-Prover-V1.5-RL xin2024deepseekprover 128 18.2 %
@@
DeepSeek-Prover-V1.5-RL + RMaxTS xin2024deepseekproverv15harnessingproofassistant 1 × 3200 21.6 %
@@
In the ProofNet benchmark, we compare REAL-Prover with several leading provers, Geodel-Prover lin2025goedel , DeepSeek-Prover-V1.5 xin2024deepseekproverv15harnessingproofassistant . For whole-proof system, the sampling budget in table means the number of generation. And for tree-search systems, the sampling budget is M×$\times$×N, where M is the total number of passes and N is the number of every step generation. The experimental results demonstrate that REAL-Prover with retrieval archieved 23.7% success rate only using supervied fine-tune. Surpassing DeepSeek-Prover-V1.5-RL under a 3200-sampling budget and DeepSeek-Prover-V1.5-RL+RMaxTS under 1$\times$3200-sampling budget, which are trained with reinforcement learning.
4.2 FATE-M Result
@@
Goedel Prover lin2025goedel 128 18.7 %
DeepSeek-Prover-V1.5-RL xin2024deepseekprover 128 31.2 %
@@
DeepSeek-Prover-V1.5-RL + RMaxTS xin2024deepseekproverv15harnessingproofassistant 64 × 64 41.8 %
@@
The performance on the FATE-M benchmark is presented in Table 2. Given the availability of open-source models and pipelines, we evaluate only Goedel-Prover and DeepSeek-Prover by using their official repositories. In this experiment, our prover achieves a state-of-the-art result with a Pass@64 of 56.7%. This demonstrates that our prover exhibits strong capability in solving college-level algebraic problems.
4.3 MiniF2F Result
The MiniF2F benchmark has been the focus of more extensive research, as summarized in Table 3. For tree search systems, serveral provers have K×$\times$×W×$\times$×N sampling budget, where K represents the number of passes, W the expansion width, and N the maximum number of proof state expansions per pass. As shown in the table, our model performs relatively poorly compared to state-of-the-art models. We consider two main factors that may contribute to this performance gap. First, MiniF2F primarily features high school-level olympiad problems, which often do not heavily rely on theorems from Mathlib. Consequently, the benefit of retrieval is diminished compared to its effectiveness on college-level problems. Second, reinforcement learning plays a significant role in the training of current state-of-the-art provers. In contrast, our model is trained solely with supervised fine-tuning, making it challenging to compete without the additional boost provided by retrieval mechanisms.
@@ 4.3 MiniF2F Result
DeepSeek-Prover-V1.5-RL xin2024deepseekprover 102400 60.2%
Leanabell-Prover-GD-RL zhang2025leanabell 128 61.1%
Goedel-Prover lin2025goedel 25600 64.7%
STP dong2502stp 25600 67.6%
@@
LLMStep welleck2023llmstep 1×32×100 27.9 %
GPT-f polu2020generativelanguagemodelingautomated 64×8×512 36.6 %
Hypertree Proof Search lample2022hypertree 64 × 5000 41.0 %
DeepSeek-Prover-V1.5-RL + RMaxTS xin2024deepseekproverv15harnessingproofassistant 32 × 6400 63.5 %
InternLM2.5-StepProver-BF wu2024internlm2 256 × 32 × 600 65.9 %
HunyuanProver v16 + BFS + DC li2024hunyuanprover 600 × 8 × 400 68.4 %
BFS-Prover xin2025bfsproverscalablebestfirsttree 2048 × 2 × 600 70.8 %
@@ 4.4 Ablation Study
In this section, we will analysis the effective of LeanSearch-PS. To compare with REAL-Prover-v1, we trained a prover model, REAL-Prover-v1-NoRet, which used the same dataset but without retrieval information. And in REAL-Prover, we do not use LeanSearch-PS for REAL-Prover-v1-NoRet model. We evaluate both models on the ProofNet and FATE-M benchmarks with results presented in Table 4. The results show that REAL-Prover-v1 achieves a performance improvement over REAL-Prover-v1-NoRet on both benchmarks. This demonstrates that the retrieval system enhances the prover’s performance on college-level mathematics problems.
We present a comparison between proofs with and without LeanSearch-PS. In Figure 3, the proof assisted by LeanSearch-PS successfully retrieves critical lemmas and completes verification, whereas the proof without LeanSearch-PS fails to verify. For problems that can be resolved both with LeanSearch-PS and without LeanSearch-PS, the one with LeanSearch-PS often leads to more accurate and human-readable proof, we provide 2 examples in Appendix E to illustrate this observation.
5 Related Work
Automated Theorem Proving in formal languages, particularly within the framework of Interactive Theorem Provers (ITPs) lean ; coq , is a rapidly advancing research area in artificial intelligence. LLM-based provers can be broadly categorized by their proof generation strategies: whole-proof generation that aims to produce complete proof in a single pass, and proof-step generation that iteratively refines proofs through ITPs.
The whole-proof generation paradigm focuses on end-to-end synthesis of formal proof scripts from theorem statements. Early attempts like DeepSeek-Prover-V1 xin2024deepseek demonstrated the feasibility of this approach in Lean 4. Subsequent improvements in DeepSeek-Prover-V1.5 xin2024deepseekproverv15harnessingproofassistant introduced hybrid techniques combining Reinforcement Learning from Proof Assistant Feedback (RLPAF) with Group Relative Policy Optimization shao2024deepseekmath , along with a ‘truncate-and-resume’ mechanism that allows error correction through partial stepwise verification. Goedel-Prover lin2025goedel addressed data scarcity through auto-formalization, first converting natural language mathematics into formal statements before generating proofs through iterative training. While these approaches benefit from large-scale supervised fine-tuning and reduced ITP interaction, they struggle with error accumulation in long logical chains and implicit state inference.
@@ 5 Related Work
This limitation has driven the development of interactive proof-step generation approaches that leverage ITPs feedback during the proving process. Systems like BFS-Prover xin2025bfsproverscalablebestfirsttree employ expert iteration with Best-First Search, using Lean compiler feedback to guide tactic generation. HunyuanProver li2024hunyuanprover enhances search efficiency through specialized critic models that evaluate intermediate proof states. The LeanDojo framework introduces ReProver yang2023leandojotheoremprovingretrievalaugmented , which adopts a distinct retrieval-based approach using Dense Passage Retriever karpukhin2020dense to directly predict relevant premises. LeanAgent kumarappan2024leanagent extends this paradigm through lifelong learning, maintaining a dynamic knowledge base to progressively master theorem proving across multiple repositories. These interactive methods typically combine expert iteration, enabling real-time adaptation to proof state changes while managing computational complexity through guided tree search.
6 Conclusions
In this work, we propose REAL-Prover, a stepwise proof system designed to tackle challenging mathematical problems. To support the development of this system, we introduce the HERALD-AF pipeline, which translates informal mathematical statements into formal Lean 4 statements, to obtain a large-scale formal dataset. We also developed a new Lean 4 interactive environment, Jixia-Interactive, which facilitates both the training of our prover model, REAL-Prover-v1, and formal proof generation. To further enhance the prover’s capabilities, we present LeanSearch-PS, a theorem retrieval system that boosts performance on the college-level mathematics problem benchmark. Additionally, to further evaluate the prover’s effectiveness, we introduce the FATE-M benchmark, where REAL-Prover achieves a state-of-the-art success rate of 56.7%.
Limitations
Despite its strong results, REAL-Prover still has two notable limitations. First, the current model is trained purely with supervised objectives. We refrained from adding reinforcement-learning (RL) stages because they demand substantially more compute and careful reward-design to avoid reward hacking. Given the evidence that RL can expose latent capabilities in large language models, incorporating RL objectives (e.g. GRPOshao2024deepseekmath ) is a promising avenue for unlocking further gains. Secondly, REAL-Prover emits only the final Lean tactics, whereas several leading provers first generate a natural-language chain of thought (CoT) and then translate it into formal steps, enabling powerful test-time reasoning amplification. Introducing an informal-reasoning layer would let the model draft, filter, and organize candidate arguments before committing to formal tactics, potentially improving both premise selection and high-level proof planning. Addressing these two aspects—RL fine-tuning and CoT-based reasoning—forms the core of our future work.
@@ Appendix A Prompt for REAL-Prover
[…]
@@ Appendix C Training Details
REAL-Prover training is performed on 8$\times$×A800 GPUs (80GB each) for approximately 18 hours, with hyperparameters detailed in Table 5.
@@
LeanSearch-PS training is performed on 4$\times$×L40 GPUs (40GB each) for approximately 12 hours with hyperparameters detailed in Table 6.
@@ Appendix E Case study
We provide 2 cases in FATE-M to illustrate that even if both REAL-Prover-v1 with LeanSearch-PS and REAL-Prover-v1-NoRet without Leansearch-PS produce the valid results, the one with premise selection often produce faster and readable proofs, as illustrated in Figure 4 and Figure 5.
@@
Figure 4: Compare the proofs with and without LeanSearch-PS. The prover in (a) uses the existing instance ‘IsAlgClosed.instInfinite’ from Mathlib, resulting in a more readable proof.
@@
(a) Proof with LeanSearch-PS.
@@
(b) Proof without LeanSearch-PS.
Figure 5: Compare the proofs with and without LeanSearch-PS. The prover in (a) uses the existing instance ‘IsPGroup.of_surjective’ from Mathlib, resulting in a more readable proof.

### v3

@@ 3.2 Benchmark
As our prover highlights its capabilities on college-level mathematics, we created our own benchmark, FATE-M, which focuses specifically on algebraic formal statements to test model’s performance on rich algebra structures. In addition, we also follow prior work [xin2024deepseekprover, theoremllama, li2024hunyuanprover, lin2025goedel] to evaluate our prover’s general capabilities using the ProofNet and MiniF2F benchmarks.
@@
The benchmark was created by extracting problems from 12 university-level abstract algebra textbooks (for a full list, see Appendix˜B), which were then formalized by students. Each formalized statement was rigorously verified by Mathlib contributors and PhD students in algebra to ensure it accurately represents the original mathematical meaning. Additionally, Lean-checked proofs are provided for all statements to guarantee their correctness.
@@
example {G : Type*} {p : ℕ} [Group G] [Fintype G] ($p$ : p.Prime) (ord : Fintype.card G = p ^ 2) : P : Subgroup G, (P.Normal) (Fintype.card P = p) :=
@@ 4.2 FATE-M Result
Goedel Prover [lin2025goedel] 32 15.6 %
DeepSeek-Prover-V1.5-SFT [xin2024deepseekprover] 128 15.9 %
DeepSeek-Prover-V1.5-RL [xin2024deepseekprover] 128 18.2 %
@@
DeepSeek-Prover-V1.5-RL + RMaxTS [xin2024deepseekproverv15harnessingproofassistant] 1 × 3200 21.6 %
@@
The performance on the FATE-M benchmark is presented in Table˜2. Given the availability of open-source models and pipelines, we evaluate only Goedel-Prover and DeepSeek-Prover by using their official repositories. In this experiment, our prover achieves a state-of-the-art result with a Pass@64 of 56.7%. This demonstrates that our prover exhibits strong capability in solving college-level algebraic problems.
4.3 MiniF2F Result
[…]
@@ 4.3 MiniF2F Result
Goedel Prover [lin2025goedel] 128 18.7 %
DeepSeek-Prover-V1.5-RL [xin2024deepseekprover] 128 31.2 %
@@
DeepSeek-Prover-V1.5-RL + RMaxTS [xin2024deepseekproverv15harnessingproofassistant] 64 × 64 41.8 %
@@
DeepSeek-Prover-V1.5-RL [xin2024deepseekprover] 102400 60.2%
Leanabell-Prover-GD-RL [zhang2025leanabell] 128 61.1%
Goedel-Prover [lin2025goedel] 25600 64.7%
STP [dong2502stp] 25600 67.6%
@@
LLMStep [welleck2023llmstep] 1×32×100 27.9 %
GPT-f [polu2020generativelanguagemodelingautomated] 64×8×512 36.6 %
Hypertree Proof Search [lample2022hypertree] 64 × 5000 41.0 %
DeepSeek-Prover-V1.5-RL + RMaxTS [xin2024deepseekproverv15harnessingproofassistant] 32 × 6400 63.5 %
InternLM2.5-StepProver-BF [wu2024internlm2] 256 × 32 × 600 65.9 %
HunyuanProver v16 + BFS + DC [li2024hunyuanprover] 600 × 8 × 400 68.4 %
BFS-Prover [xin2025bfsproverscalablebestfirsttree] 2048 × 2 × 600 70.8 %
@@ 4.4 Ablation Study
In this section, we will analysis the effective of LeanSearch-PS. To compare with REAL-Prover-v1, we trained a prover model, REAL-Prover-v1-NoRet, which used the same dataset but without retrieval information. And in REAL-Prover, we do not use LeanSearch-PS for REAL-Prover-v1-NoRet model. We evaluate both models on the ProofNet and FATE-M benchmarks with results presented in Table˜4. The results show that REAL-Prover-v1 achieves a performance improvement over REAL-Prover-v1-NoRet on both benchmarks. This demonstrates that the retrieval system enhances the prover’s performance on college-level mathematics problems.
We present a comparison between proofs with and without LeanSearch-PS. In Figure˜3, the proof assisted by LeanSearch-PS successfully retrieves critical lemmas and completes verification, whereas the proof without LeanSearch-PS fails to verify. For problems that can be resolved both with LeanSearch-PS and without LeanSearch-PS, the one with LeanSearch-PS often leads to more accurate and human-readable proof, we provide 2 examples in Appendix˜E to illustrate this observation.
@@
5 Conclusion and Future Work
In this work, we propose REAL-Prover, a stepwise proof system designed to tackle challenging mathematical problems. To support its development, we introduce the HERALD-AF pipeline, which translates informal mathematical statements into formal Lean 4 representations, enabling the construction of a large-scale formal dataset. We also develop a new interactive Lean 4 environment, Jixia-Interactive, which facilitates both the training of our prover model, REAL-Prover-v1, and efficient formal proof generation. Furthermore, we present LeanSearch-PS, a theorem retrieval system that improves theorem selection and boosts overall proof success on college-level mathematical benchmarks. To comprehensively evaluate the prover’s effectiveness, we propose the FATE-M benchmark, on which REAL-Prover achieves a state-of-the-art success rate of 56.7%.
Despite these promising results, several limitations remain. The current model is trained purely with supervised objectives, and we have not yet incorporated reinforcement learning (RL) due to its computational demands. Nevertheless, prior studies suggest that RL can unlock latent reasoning capabilities in large language models, indicating that RL-based fine-tuning (e.g., GRPO [shao2024deepseekmath]) constitutes a promising avenue for enhancing performance. In addition, REAL-Prover emits only the final Lean tactics, whereas several leading provers first generate a natural-language chain of thought (CoT) and then translate it into formal steps, enabling powerful test-time reasoning amplification. Introducing an informal-reasoning layer would let the model draft, filter, and organize candidate arguments before committing to formal tactics, potentially improving both premise selection and high-level proof planning.
Future work will focus on these two complementary directions: incorporating RL-based fine-tuning to deepen reasoning capabilities and generalization, and integrating CoT-driven informal reasoning to bridge natural and formal mathematical discourse. These extensions are expected to yield a more autonomous, interpretable, and robust formal reasoning system, advancing the broader objective of scalable, human-aligned mathematical intelligence.
@@ Appendix A Prompt for REAL-Prover
In Lean, a formal proof is a fully constructed proof term that is type-checked and verified by the kernel. It represents a complete and correct derivation of a proposition. The state after tactics refers to the intermediate proof state during tactic-based proof construction. It includes the list of remaining goals and the local context at that point. Relationship: 1. Tactics are procedural tools used to incrementally construct a formal proof. 2. Each tactic transforms the current proof state by solving or reducing goals. 3. The state after a tactic reflects the goals that still need to be proven after that tactic has been applied. 4. Once all goals are solved, Lean assembles the underlying proof terms generated by the tactics into a complete formal proof. 5. This final term is then type-checked by the kernel to ensure correctness. In essence, the state after tactics shows where you are in the process of building a formal proof — it’s a snapshot of what’s left to do before the proof is complete. Here is the FORMAL PROOF before the current state: import Mathlib open BigOperators open Real Nat Topology theorem parallelogram_midpoints_coincide_extracted {A B C D : ℂ} : (A + B + C + D) / 4 = (A + C) / 2 (B + D) / 2 = (A + C) / 2 B.re + D.re = 2 * ((A + C) / 2).re := by simp [add_comm, add_left_comm, add_assoc] simp [div_eq_mul_inv, mul_add, mul_comm, mul_left_comm] intro h h Here is the current STATE: A : B : C : D : h : (A + (B + (C + D))) * 4⁻¹ = (A + C) * 2⁻¹ h : B + D = A + C B.re + D.re = A.re + C.re Please generate a tactic in lean4 to solve the state. And here’re some theorems that may be helpful: ID:0 Formal name:Complex.add_re Informal name: Real Part of Sum of Complex Numbers Formal statement: (z w : ℂ), (z + w).re = z.re + w.re ID:5 Formal name: add_add_add_comm Informal name: Commutativity and Associativity of Addition in Additive Commutative Semigroups Formal statement: {G : Type u_3} [inst : AddCommSemigroup G] (a b c d : G), a + b + (c + d) = a + c + (b + d)
@@ Appendix C Training Details
REAL-Prover training is performed on $\timesA8$00 GPUs (80GB each) for approximately 18 hours, with hyperparameters detailed in Table˜5.
@@
LeanSearch-PS training is performed on $\timesL4$0 GPUs (40GB each) for approximately 12 hours with hyperparameters detailed in Table˜6.
@@ Appendix E Case study
We provide 2 cases in FATE-M to illustrate that even if both REAL-Prover-v1 with LeanSearch-PS and REAL-Prover-v1-NoRet without Leansearch-PS produce the valid results, the one with premise selection often produce faster and readable proofs, as illustrated in Appendix˜E and Appendix˜E.
@@
(c) Compare the proofs with and without LeanSearch-PS. The prover in (a) uses the existing instance ‘IsAlgClosed.instInfinite’ from Mathlib, resulting in a more readable proof.
@@
(d) Proof with LeanSearch-PS.
@@
(e) Proof without LeanSearch-PS.
(f) Compare the proofs with and without LeanSearch-PS. The prover in (a) uses the existing instance ‘IsPGroup.of_surjective’ from Mathlib, resulting in a more readable proof.